An open at both ends pipe has a length of 50cm. What are three lowest

frequencies to which it will resonate? v=340m/s]



1
Expert's answer
2021-02-05T14:06:23-0500

"f=\frac{nv}{2L}"


"f_1=\frac{1\cdot 340}{2\cdot 0.5}=340\ (Hz)"


"f_2=\frac{2\cdot 340}{2\cdot 0.5}=680\ (Hz)"


"f_3=\frac{3\cdot 340}{2\cdot 0.5}=1020\ (Hz)"